# Realistic Cost to Execute Practical Quantum Circuits using Direct Clifford+T Lattice Surgery Compilation

### v2

@@ Abstract.
In this article, we report a resource estimation pipeline that explicitly compiles quantum circuits expressed using the Clifford+T gate set into a surface code lattice surgery instruction set. The cadence of magic state requests from the compiled circuit enables the optimization of magic state distillation and storage requirements in a post-hoc analysis. To compile logical circuits into lattice surgery operations, we build upon the open-source Lattice Surgery Compiler. The revised compiler operates in two stages: the first translates logical gates into an abstract, layout-independent instruction set; the second compiles these into local lattice surgery instructions that are allocated to hardware tiles according to a specified resource layout. The second stage retains logical parallelism while avoiding resource contention in the fault-tolerant layer, aiding realism. Additionally, users can specify dedicated tiles at which magic states are replenished, enabling resource costs from the logical computation to be considered independently from magic state distillation and storage. We demonstrate the applicability of our pipeline to large, practical quantum circuits by providing resource estimates for the ground state estimation of molecules. We find that, unless carefully considered, the resource costs of magic state storage can dominate in real circuits which have variable magic state consumption rates.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
In quantum computing, promises of exponential speed-ups for certain classes of computationally hard problems have spurred an intense focus on research and development of systems capable of fulfilling these promises (Gibney, 2019; Parker et al., 2022). It is generally understood that, due to the fragility of quantum states, it is necessary to encode logical qubit states within the joint Hilbert space of many physical qubits using what are called quantum error-correcting codes (QECC) (Gottesman, 1997; Terhal, 2015; Campbell et al., 2017). Currently, the surface code is the leading QECC due to the locality of its error checks (implying straightforward hardware implementation) and high error threshold (implying greater fault-tolerance) (Kitaev, 2003; Fowler et al., 2012; Litinski, 2019a). While it can be said that we are now in the fault-tolerant era due to there having been demonstrations of fault-tolerance by major hardware providers (Ryan-Anderson et al., 2021; Abobeih et al., 2022; goo, 2023), the overhead of surface code quantum computing is large enough that leading systems may not have enough qubits to implement practical quantum algorithms fault-tolerantly for a long time. Still, there has already been much discussion and refinement in the literature regarding strategies to compile logical circuits into fault-tolerant operations on the surface code [^1] We will call this concept fault-tolerant compilation., with lattice surgery offering the leading instruction set (Horsman et al., 2012; Herr et al., 2017a, b; Brown et al., 2017; Fowler and Gidney, 2018; Litinski and von Oppen, 2018; Litinski, 2019a; Beverland et al., 2022a, b; Chamberland and Campbell, 2022b, a), and with some open-source compilers already having been developed (Paler and Fowler, 2020; Watkins et al., 2023). The purpose of this research direction is twofold. First, and most obviously, research into fault-tolerant compilation and the development of fault-tolerant compilers serves to prepare us to execute quantum algorithms fault-tolerantly on future quantum computers. Secondly, and more immediately, these efforts aid us in estimating resource costs for and evaluating the potential utility of future quantum computers, which is important for informing current directions in quantum computing research.
@@
With this in mind, we argued in Ref. (LeBlond et al., 2023b) for the importance of developing extensible fault-tolerant compilation platforms with verification. One purpose for doing so is that the resource overheads for different fault-tolerant implementations (e.g. using different instruction sets or even QECC) of quantum algorithms could then be readily compared, aiding in the optimization of the fault-tolerant layer of the stack. The open-source Lattice Surgery Compiler (LSC) offers a starting point in this direction for surface code lattice surgery specifically (Watkins et al., 2023). The LSC was originally designed for sequential Pauli-based computation, which is a surface code compilation framework that requires the global re-write of a logical circuit into sequential Pauli product measurements with support over the whole qubit register (Litinski, 2019a). However, it also supports direct Clifford+T compilation using techniques from Refs. (Fowler et al., 2012; Horsman et al., 2012; Litinski and von Oppen, 2018; Litinski, 2019a). Direct Clifford+T compilation is sometimes considered preferable to sequential Pauli-based computation due to its ability to partially preserve parallelism from the logical circuit in the fault-tolerant layer (Beverland et al., 2022a). In this work, we have taken advantage of the extensibility of the LSC to directly compile Clifford+T circuits into a lattice surgery instruction set similar to the one proposed in Ref. (Beverland et al., 2022a), using its Bell-based methodology to compile long-range gates. We have chosen to implement a local lattice surgery-based surface code instruction set because of its easy reduction into a limited set of primitive hardware circuits acting independently on one or two hardware tiles that can be readily verified (LeBlond et al., 2023a), paving the way toward a verified end-to-end compilation pipeline for quantum circuits. While this choice adds roughly a factor of 2 overhead to some operations, we suspect that this is outweighed by a greater ease in decoding and instruction scheduling in the physical qubit layer compared with the use of extended ancillae to mediate long-range interactions.
@@
The LSC is not only extensible to different instruction sets but also to methods for routing instructions and exploiting their logical parallelism. In the LSC, a compilation pipeline refers to the complete process of compiling an input circuit to output lattice surgery instructions, including translations to any intermediate representations. In our upgraded LSC, we have implemented a new compilation pipeline (called the wave pipeline) that efficiently detects circuit-level parallelism and combines this information with hardware layout constraints to map logical circuits into time-slices of parallelizable fault-tolerant operations. By effectively handling contention for space by logically parallel operations, the wave pipeline allows the LSC to find a realistic number of logical time-slices (as well as a realistic active volume [^2] The active volume is the active portion of the total space-time volume of the circuit including both the cost of logical operations and the cost of idling logical qubits, since logical errors can accumulate in both. Note that this definition slightly differs from the one introduced in Ref. (Litinski and Nickerson, 2022).) required to execute a logical circuit fault-tolerantly. Importantly, the wave pipeline implies a change in compilation philosophy as compared with the streamed pipeline from the original LSC that was described in Ref. (Watkins et al., 2023). Namely, the wave pipeline assumes that compilation takes place ahead of circuit execution while the streamed pipeline assumes that compilation takes place in tandem with circuit execution. We will discuss the implications of this choice throughout the paper, but the main benefits of the former approach are that it enables greater exploitation of circuit-level parallelism as well as the post-hoc optimization of resources dedicated to magic state distillation and storage, with drawbacks including potential limitations on circuit size due to classical memory constraints. We believe that future work should combine these approaches.
@@
Though many recent efforts towards the resource estimation of practical quantum circuits do consider the costs of fault-tolerance and are often quite sophisticated, they typically stop short of explicit fault-tolerant compilation (Fowler et al., 2012; Litinski, 2019a; Gouzien and Sangouard, 2021; Chamberland et al., 2022; Chamberland and Campbell, 2022b; Beverland et al., 2022b; Kim et al., 2022). Ref. (Beverland et al., 2022b) introduced a particularly flexible resource estimator that nevertheless makes several important concessions, such as assuming a constant rate of magic state consumption, using a worst-case formula for synthesis of single-qubit rotations, and neglecting that unused tiles will not contribute to the logical error rate. We believe that explicit compilation within the fault-tolerant layer is necessary to accurately assess the impact of such factors on resource requirements. In this paper, we combine elements of the framework from (Beverland et al., 2022b) with the output of our revised LSC to estimate resources for large, practical quantum circuits. To this end, we explicitly synthesize single-qubit rotations and use the active volume of the final (compiled) circuit to calculate the total logical error rate. We also present a novel method to estimate the resource requirements for magic state distillation and storage. First, the cadence of magic state requests from the compiled output is used to determine the minimum number of magic state factories needed to keep up with magic state consumption in the circuit. Then, knowing the magic state production and consumption rates, the number of magic states in storage after each distillation cycle can be tracked; this turns out to figure significantly in resource costs for the algorithm. We calculate contributions to space, time, and active volume from the logical, distillation, and storage portions of the quantum computation and optimize the total of each (or any) of these resources over the code distances under the constraint that the total error rate lies beneath a threshold. As a demonstration of our methodology, we present resource estimates for one Trotter step of a novel ground-state estimation circuit (Kornell and Selinger, 2023). We find that, unless care is taken in optimizing the usage of resources dedicated to magic state distillation, the costs of magic state storage can quickly dominate; using a minimal number of continuously operating magic state factories that meets the consumption requirements of the circuit is not sufficient. Though we consider an abstract model of quantum computing hardware in this paper, we show that the output of our resource analysis (expressed in terms of resources from the fault-tolerant layer such as hardware tiles and logical time-slices) can be readily combined with hardware-specific resource analyses to produce resource costs in terms of space and time [^3] The resource estimation pipeline considered here was designed to be compatible with the trapped-ion surface code compiler (TISCC) (LeBlond et al., 2023a)..
@@ 2. Lattice Surgery Compilation
As mentioned in Sec. 1, the primary goal of this work is to enable accurate estimation of the resources required to run large quantum circuits on an abstract model of quantum computing hardware. For a high-level overview of our methodology, see Fig. 1. Explicit lattice surgery compilation within the logical block [^4] The logical block is the portion of the quantum computer dedicated to executing the logical circuit and excludes resources dedicated to magic state distillation and storage. yields total estimates of resources such as logical time-slices, hardware tiles, and active volume within the fault-tolerant layer [^5] For lack of a better term, we will call these abstract resources fault-tolerant resources. that figure into our full resource estimation methodology, which will be presented in Sec. 4. The purpose of Sec. 2 is to provide details for our approach to lattice surgery compilation, while the costs of generating and managing magic states, as well as the method for integrating these with logical execution costs, will be reserved for Sec. 4. First, in Sec. 2.1, we will share some general considerations about lattice surgery compilation. In Sec. 2.2 we outline our strategy for lattice surgery compilation, and in Sec. 2.3 we discuss the LSC, the open-source software package that we have built upon, as well as some specific ways in which we have modified the LSC to suit our compilation strategy.
@@ 2.1. General Considerations
Multiple layers of complexity become apparent while considering fault-tolerant compilation (LeBlond et al., 2023b). Layers of the stack can be coarse-grained into the logical layer, in which a quantum circuit is expressed by logical gates acting on logical qubits; the fault-tolerant layer, in which logical gates have been re-expressed in terms of fault-tolerant operations on a QECC; and the hardware layer, in which fault-tolerant operations are expressed in terms of native hardware operations. However, these distinctions are not concrete. Not only can these layers of the stack be broken down further, choices within one layer may be co-dependent with choices made in other layers. To take one example, the specifics of how to implement a logical instruction within the fault-tolerant layer depend on the layout of resources on the underlying hardware architecture, which is typically seen as an input to fault-tolerant compilation (see Fig. 1), but may also be influenced by hardware constraints such as the need to divide resources across multiple processors. As we will see, even the gate set used in the logical layer can depend on one’s preferred set of fault-tolerant operations, which in turn can depend on hardware-layer constraints.
@@
For surface codes, there has already been much discussion in the literature about compilation strategies; Refs. (Horsman et al., 2012; Herr et al., 2017a, b; Brown et al., 2017; Fowler and Gidney, 2018; Litinski and von Oppen, 2018; Litinski, 2019a; Beverland et al., 2022a, b; Chamberland and Campbell, 2022b, a) represent a sub-set of the literature pertaining to surface code compilation using lattice surgery specifically. We will not comprehensively review this literature here, but will briefly discuss a few nuances which serve to illustrate our points above. To provide a setting in which to perform lattice surgery compilation, the arrangement of resources on the hardware is often specified in terms of an abstract layout of tiles on a grid. These tiles are an abstraction of the area of hardware required to support a single surface code patch at the minimum code distances required by the computation [^6] See Ref. (LeBlond et al., 2023a) for the more precise definition of logical tile that is compatible with this work.. Given a layout specification, the implementation of a logical instruction within the fault-tolerant layer can depend on the availability and locations of surface code patches encoding resource states and/or the availability of connected paths of free tiles (routes) between surface code patches encoding logical data. Once a resource state and/or route is found for a given instruction, one or several ancilla patches may need to be prepared to facilitate the interaction [^7] Surface code compilation techniques, especially those that use lattice surgery, typically rely on the preparation of surface code patches encoding ancillae that were not present in the original logical circuit. The spatial overhead for ancillae can be significant (Litinski, 2019a; Beverland et al., 2022a; Chamberland and Campbell, 2022b).. Therefore, the compilation of logical gates into fault-tolerant operations on the surface code is often broken into two stages, where in the first stage the logical gate is compiled into elements of an abstract, layout-independent instruction set while in the second stage the availabilities of resource states and routes are accounted for to implement these abstract instructions in terms of instructions acting on specific hardware tiles (Watkins et al., 2023). While it has been shown that layout optimization in lattice surgery compilation is NP hard given a particular compilation strategy (Herr et al., 2017a, b), authors frequently hand-optimize layouts to better suit their choice of abstract instruction set within the fault-tolerant layer so that the overhead resulting from routing instructions in the second compilation stage is minimized [^8] See Refs. (Litinski, 2019a) and (Litinski, 2019b) for vivid examples of hardware layouts being hand-designed to suit a lattice surgery compilation scheme while exploring various space-time trade-offs..
@@
In some strategies, such as those presented in Refs. (Litinski, 2019a; Chamberland and Campbell, 2022b; Beverland et al., 2022b), instead of compiling logical circuits into fault-tolerant operations gate-by-gate, a global re-write of the original logical circuit first eliminates Clifford gates, resulting in a circuit that primarily contains the Pauli product measurements that are native to surface code lattice surgery (Fowler and Gidney, 2018). The details of this circuit re-write can depend on one’s choice of abstract instruction set within the fault-tolerant layer (in this case, the set of allowed Pauli product measurements) as well as the hardware layout (Litinski, 2019a; Beverland et al., 2022b; Chamberland and Campbell, 2022b). Furthermore, the set of allowed Pauli product measurements can depend on the relative ease with which one can implement twist defects on their underlying hardware, which influences whether the logical layer is re-written to include Pauli product measurements with Y Paulis (Chamberland and Campbell, 2022a, b). Therefore, the hardware layer of the compilation stack has implications that propagate all the way up to the logical layer. As mentioned in Sec. 1, direct Clifford+T compilation is sometimes preferred to the Clifford elimination strategy (sometimes called sequential Pauli-based computation) due to its ability to preserve logical circuit-level parallelism. Within this context, there are also many choices e.g. whether to implement a CNOT gate using extended ancilla patches as in Ref. (Litinski and von Oppen, 2018) or using long-range Bell states as in Ref. (Beverland et al., 2022a), whether to implement an S gate using teleportation as in Ref. (Beverland et al., 2022a) or by braiding twist defects as in Ref. (Brown et al., 2017), or in how to perform a patch rotation following a transversal Hadamard gate (see Refs. (Vuillot et al., 2019; Litinski, 2019a; Beverland et al., 2022a; Horsman et al., 2012; Fowler and Gidney, 2018) for different protocols); the optimal choice will be similarly interdependent with choices of layout and hardware constraints.
@@
These considerations have motivated the realization that the three mentioned layers of the stack are highly inter-dependent and their co-optimization is formidable, if not impossible. At the same time, because the fault-tolerant layer is expected to dramatically increase the resource overhead of quantum computing, any end-to-end resource estimation pipeline for quantum circuits ought to include reasonable choices for it. To summarize this section, we take note that choices must be made in (at least) the following places: (a) a logical gate set with which to express input circuits (e.g. Clifford+T or Pauli product rotations), (b) a lattice surgery-native logical instruction set to translate input circuits to (e.g. the allowed set of Pauli product measurements), (c) an abstract layout specifying regions on the hardware dedicated to storing logical qubits, routing long-range instructions, and generating resource states (d) a protocol for implementing the lattice surgery-native logical instruction set using the layout (including methods for scheduling and routing these instructions as well as generating and managing any required resource states), and (e) a protocol for further compiling routed and scheduled lattice surgery instructions into native hardware gates. As we are convinced that extensible compilation and resource estimation platforms would greatly aid the community in optimizing resources within the fault-tolerant layer and culling the spectrum of propositions for surface code compilation strategies, we present one set of choices in this paper and expand the functionality of the LSC to suit.

### v4

@@ Abstract.
We report a resource estimation pipeline that explicitly compiles quantum circuits expressed using the Clifford+T gate set into a surface code lattice surgery instruction set. The cadence of magic state requests from the compiled circuit enables the optimization of magic state distillation and storage requirements in a post-hoc analysis. To compile logical circuits into lattice surgery operations, we build upon the open-source Lattice Surgery Compiler. The revised compiler operates in two stages: the first translates logical gates into an abstract, layout-independent instruction set; the second compiles these into local lattice surgery instructions that are allocated to hardware tiles according to a specified resource layout. The second stage retains logical parallelism while avoiding resource contention in the fault-tolerant layer, aiding realism. Additionally, users can specify dedicated tiles at which magic states are replenished, enabling resource costs from the logical computation to be considered independently from magic state distillation and storage. We demonstrate the applicability of our pipeline to large, practical quantum circuits by providing resource estimates for the ground state estimation of molecules. We find that variable magic state consumption rates in real circuits can cause the resource costs of magic state storage to dominate unless production is varied to suit.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXcopyright: licensedusgovmixedjournalyear: 2024journalvolume: 5journalnumber: 4article: 25publicationmonth: 10doi: 10.1145/3689826
@@ 1. Introduction
In quantum computing, promises of exponential speed-ups for certain classes of computationally hard problems have spurred an intense focus on research and development of systems capable of fulfilling these promises (Gibney, 2019; Parker et al., 2022). It is generally understood that, due to the fragility of quantum states, it is necessary to encode logical qubit states within the joint Hilbert space of many physical qubits using what are called quantum error-correcting codes (QECC) (Gottesman, 1997; Terhal, 2015; Campbell et al., 2017). Currently, the surface code is the leading QECC due to the locality of its error checks (implying straightforward hardware implementation) and high error threshold (implying greater fault-tolerance) (Kitaev, 2003; Fowler et al., 2012; Litinski, 2019a). While it can be said that we are near the fault-tolerant era because there have been demonstrations of fault-tolerance by major hardware providers (Ryan-Anderson et al., 2021; Abobeih et al., 2022; goo, 2023; da Silva et al., 2024), the overhead of surface code quantum computing is large enough that leading systems may not have enough qubits to implement practical quantum algorithms fault-tolerantly for a long time. Still, there has already been much discussion and refinement in the literature regarding strategies to compile logical circuits into fault-tolerant operations on the surface code [^1] We will call this concept fault-tolerant compilation., with lattice surgery offering the leading instruction set (Horsman et al., 2012; Herr et al., 2017a, b; Brown et al., 2017; Fowler and Gidney, 2018; Litinski and von Oppen, 2018; Litinski, 2019a; Beverland et al., 2022a, b; Chamberland and Campbell, 2022b, a), and with some open-source compilers already having been developed (Paler and Fowler, 2020; Watkins et al., 2024). The purpose of this research direction is twofold. First, and most obviously, research into fault-tolerant compilation and the development of fault-tolerant compilers serves to prepare us to execute quantum algorithms fault-tolerantly on future quantum computers. Secondly, and more immediately, these efforts aid us in estimating resource costs for and evaluating the potential utility of future quantum computers, which is important for informing current directions in quantum computing research.
@@
With this in mind, we argued in Ref. (LeBlond et al., 2023b) for the importance of developing extensible fault-tolerant compilation platforms with verification. One purpose for doing so is that the resource overheads for different fault-tolerant implementations (e.g. using different instruction sets or even QECC) of quantum algorithms could then be compared, aiding in the optimization of the fault-tolerant layer of the stack. The open-source Lattice Surgery Compiler (LSC) offers a starting point in this direction for surface code lattice surgery specifically (Watkins et al., 2024). The LSC was originally designed for sequential Pauli-based computation, which is a surface code compilation framework that requires the global re-write of a logical circuit into sequential Pauli product measurements with support over the whole qubit register (Litinski, 2019a). However, it also supports direct Clifford+T compilation using techniques from Refs. (Fowler et al., 2012; Horsman et al., 2012; Litinski and von Oppen, 2018; Litinski, 2019a). Direct Clifford+T compilation is sometimes considered preferable to sequential Pauli-based computation due to its ability to partially preserve parallelism from the logical circuit in the fault-tolerant layer (Beverland et al., 2022a). In this work, we have taken advantage of the extensibility of the LSC to directly compile Clifford+T circuits into a lattice surgery instruction set similar to the one proposed in Ref. (Beverland et al., 2022a), using its Bell-based methodology to compile long-range gates. We have chosen to implement a local lattice surgery-based surface code instruction set because of its easy reduction into a limited set of primitive hardware circuits acting independently on one or two hardware tiles that can be readily verified (LeBlond et al., 2023a), paving the way toward a verified end-to-end compilation pipeline for quantum circuits. While this choice adds roughly a factor of 2 overhead to some operations, we suspect that this is outweighed by a greater ease in decoding and instruction scheduling in the physical qubit layer compared with the use of extended ancillae to mediate long-range interactions.
@@
The LSC is not only extensible to different instruction sets but also to methods for routing instructions and exploiting their logical parallelism. In the LSC, a compilation pipeline refers to the complete process of compiling an input circuit to output lattice surgery instructions, including translations to any intermediate representations. In our upgraded LSC, we have implemented a new compilation pipeline (called the wave pipeline) that efficiently detects circuit-level parallelism and combines this information with hardware layout constraints to map logical circuits into time-slices of parallelizable fault-tolerant operations. By effectively handling contention for space by logically parallel operations, the wave pipeline allows the LSC to find a realistic number of logical time-slices (as well as a realistic active volume [^2] The active volume is the active portion of the total space-time volume of the circuit including both the cost of logical operations and the cost of idling logical qubits, since logical errors can accumulate in both. Note that this definition slightly differs from the one introduced in Ref. (Litinski and Nickerson, 2022).) required to execute a logical circuit fault-tolerantly. Importantly, the wave pipeline implies a change in compilation philosophy as compared with the streamed pipeline from the original LSC that was described in Ref. (Watkins et al., 2024). Namely, the wave pipeline assumes that compilation takes place ahead of circuit execution while the streamed pipeline assumes that compilation takes place in tandem with circuit execution. We will discuss the implications of this choice throughout the paper, but the main benefits of the former approach are that it enables greater exploitation of circuit-level parallelism as well as the post-hoc optimization of resources dedicated to magic state distillation and storage, with drawbacks including potential limitations on circuit size due to classical memory constraints. We believe that future work should combine these approaches.
@@
Though many recent efforts towards the resource estimation of practical quantum circuits do consider the costs of fault-tolerance and are often quite sophisticated, they typically stop short of explicit fault-tolerant compilation (Fowler et al., 2012; Litinski, 2019a; Gouzien and Sangouard, 2021; Chamberland et al., 2022; Chamberland and Campbell, 2022b; Beverland et al., 2022b; Kim et al., 2022). Ref. (Beverland et al., 2022b) introduced a particularly flexible resource estimator that nevertheless makes several important concessions, such as assuming a constant rate of magic state consumption, using a worst-case formula for synthesis of single-qubit rotations, and neglecting that unused tiles will not contribute to the logical error rate. We believe that explicit compilation within the fault-tolerant layer is necessary to accurately assess the impact of such factors on resource requirements. In this paper, we combine elements of the framework from (Beverland et al., 2022b) with the output of our revised LSC to estimate resources for large, practical quantum circuits. To this end, we explicitly synthesize single-qubit rotations and use the active volume of the final (compiled) circuit to calculate the total logical error rate. We also present a novel method to estimate the resource requirements for magic state distillation and storage. First, the cadence of magic state requests from the compiled output is used to determine the minimum number of magic state factories needed to keep up with magic state consumption in the circuit. Then, knowing the magic state production and consumption rates, the number of magic states in storage after each distillation cycle can be tracked. We calculate contributions to space, time, and active volume from the logical, distillation, and storage portions of the quantum computation and optimize the total of each (or any) of these resources over the code distances under the constraint that the total error rate lies beneath a threshold. As a demonstration of our methodology, we present resource estimates for one Trotter step of a novel ground-state estimation circuit (Kornell and Selinger, 2023). We find that using a minimal number of continuously operating magic state factories that meets the consumption requirements of the circuit causes the spatial cost of magic state storage to dominate, which in turn causes huge total space-time volumes. To bring storage costs down, we propose a method to match the real-time magic state production rate to the consumption rate by turning factories off and on as needed. Using this technique, we find more than three orders of magnitude improvement in space-time volume for our largest circuit. Finally, though we consider an abstract model of quantum computing hardware in this paper, we show that the output of our resource analysis (expressed in terms of resources from the fault-tolerant layer such as hardware tiles and logical time-slices) can be readily combined with hardware-specific resource analyses to produce resource costs in terms of space and time [^3] The resource estimation pipeline considered here was designed to be compatible with the trapped-ion surface code compiler (TISCC) (LeBlond et al., 2023a)., which completes an end-to-end resource estimation tool-chain.
@@
1.1. Contributions
  • We present the first end-to-end resource estimation pipeline for quantum circuits that utilizes explicit lattice surgery compilation to obtain (a) a realistic number of logical slices, (b) a realistic active volume, and (c) a setting for the optimization of magic state distillation and storage based upon “real-time” magic state consumption rates.
  • To accomplish this, we have made advances to the state-of-the-art in lattice surgery compilation tooling:
    • We improve the open-source Lattice Surgery Compiler (LSC) (Watkins et al., 2024) to detect logical parallelism and translate it into parallelism of fault-tolerant operations while handling contention for routes and resource states.
    • We make other improvements to the LSC such as improved layout generation, more flexible magic state handling, and an improved output format that facilitates its integration within our tool-chain.
    • We present a detailed performance analysis using both randomly-generated and practical circuits. We find that the scalings of both compilation and computation time are consistent with theoretical predictions and compare the performance characteristics of layouts with different ratios of data and ancilla qubits.
  • Our pipeline paves the way toward verified end-to-end (logical-to-hardware) compilation for fault-tolerant processors through implementing a local lattice surgery compilation layer that connects logical lattice instructions to the finite set of logically verified lattice surgery primitives acting on pairs of hardware tiles from Ref. (LeBlond et al., 2023a).
  • We demonstrate the usage of our pipeline by generating resource estimates for the execution of ground state estimation circuits on a hypothetical trapped-ion processor design. We also explore the distillation-storage trade space for these circuits and mitigate the large magic state storage requirements that we discover.
@@ 2. Lattice Surgery Compilation
As mentioned in Sec. 1, the primary goal of this work is to enable accurate estimation of the resources required to run large quantum circuits on an abstract model of quantum computing hardware. For a high-level overview of our methodology, see Fig. 1. The left-hand side of the diagram, which pertains to explicit lattice surgery compilation within the logical block [^4] The logical block is the portion of the quantum computer dedicated to executing the logical circuit and excludes resources dedicated to magic state distillation and storage., yields total estimates of resources from the fault-tolerant layer such as logical time-slices, hardware tiles, and active volume [^5] For lack of a better term, we will call these abstract resources fault-tolerant resources. that figure into the resource analysis scheme represented by the right-hand side of the diagram. The purpose of Sec. 2 is to provide details for our approach to lattice surgery compilation, while the costs of generating and managing magic states, as well as the method for integrating these with logical execution costs, will be reserved for Sec. 4. First, in Sec. 2.1, we will share some general considerations about lattice surgery compilation. In Sec. 2.2 we outline our strategy for lattice surgery compilation, and in Sec. 2.3 we discuss the LSC, the open-source software package that we have built upon, as well as some specific ways in which we have modified the LSC to suit our compilation strategy.
@@ 2.1. General Considerations
Multiple layers of complexity become apparent while considering fault-tolerant compilation (LeBlond et al., 2023b). Layers of the stack can be coarsely divided into the logical layer, in which a quantum circuit is expressed by logical gates acting on logical qubits; the fault-tolerant layer, in which logical gates have been re-expressed in terms of fault-tolerant operations on a QECC; and the hardware layer, in which fault-tolerant operations are expressed in terms of native hardware operations. However, these distinctions are not concrete. Not only can these layers of the stack be broken down further, choices within one layer may be co-dependent with choices made in other layers. To take one example, the specifics of how to implement a logical instruction within the fault-tolerant layer depend on the arrangement of logical resources on the underlying hardware architecture. This layout can be treated as the subject of optimization by a fault-tolerant compiler, but may also be influenced by hardware limitations such as the need to divide resources across multiple processors.
@@
For surface codes, there has already been much discussion in the literature about compilation strategies; Refs. (Horsman et al., 2012; Herr et al., 2017a, b; Brown et al., 2017; Fowler and Gidney, 2018; Litinski and von Oppen, 2018; Litinski, 2019a; Beverland et al., 2022a, b; Chamberland and Campbell, 2022b, a) represent a sub-set of the literature pertaining to surface code compilation using lattice surgery specifically. We will not comprehensively review this literature here, but will briefly discuss a few nuances which serve to illustrate our points above. To provide a setting in which to perform lattice surgery compilation, the arrangement of logical resources on the hardware is often specified in terms of an abstract layout of tiles on a grid (see Fig. 4 for examples). These tiles are an abstraction of the area of hardware required to support a single surface code patch at the minimum code distances required by the computation [^6] See Ref. (LeBlond et al., 2023a) for the more precise definition of logical tile that is compatible with this work.. Given a layout specification, the implementation of a logical instruction within the fault-tolerant layer can depend on the availability and locations of surface code patches encoding resource states and/or the availability of connected paths of free tiles (routes) between surface code patches encoding logical data. Once a resource state and/or route is found for a given instruction, one or several ancilla patches may need to be prepared to facilitate the interaction [^7] Surface code compilation techniques, especially those that use lattice surgery, typically rely on the preparation of surface code patches encoding ancillae that were not present in the original logical circuit. The spatial overhead for ancillae can be significant (Litinski, 2019a; Beverland et al., 2022a; Chamberland and Campbell, 2022b).. Therefore, the compilation of logical gates into fault-tolerant operations on the surface code is often broken into two stages. In the first stage, each logical gate is compiled into elements of an intermediate representation containing instructions that pertain to the fault-tolerant layer but are still abstracted from layout specifics. In the second stage, the availabilities of resource states and routes are accounted for to implement these abstract instructions in terms of fault-tolerant operations acting on specific hardware tiles (Watkins et al., 2024). While it has been shown that layout optimization in lattice surgery compilation can be NP hard (Herr et al., 2017a, b), authors frequently hand-optimize layouts to suit particular compilation schemes so that resource overheads resulting from routing instructions in the second compilation stage are minimized [^8] See Refs. (Litinski, 2019a) and (Litinski, 2019b) for vivid examples of hardware layouts being hand-designed to suit a lattice surgery compilation scheme while exploring various space-time trade-offs..
@@
In some strategies, such as those presented in Refs. (Litinski, 2019a; Chamberland and Campbell, 2022b; Beverland et al., 2022b), instead of compiling logical circuits into fault-tolerant operations gate-by-gate, a global re-write of the original logical circuit first eliminates Clifford gates, resulting in a circuit that can be reduced to Pauli product measurements with support over the whole qubit register (these operations are native to surface code lattice surgery (Fowler and Gidney, 2018)). The details of this circuit re-write can depend on one’s choice of allowed Pauli product measurements as well as the hardware layout (Litinski, 2019a; Beverland et al., 2022b; Chamberland and Campbell, 2022b). Furthermore, the set of allowed Pauli product measurements can depend on the relative ease with which one can implement twist defects on their underlying hardware, which influences whether the logical layer is re-written to include Pauli product measurements with Y Paulis (Chamberland and Campbell, 2022a, b). Therefore, the hardware layer of the compilation stack has implications that propagate all the way up to the logical layer. As mentioned in Sec. 1, direct Clifford+T compilation is sometimes preferred to the Clifford elimination strategy (sometimes called sequential Pauli-based computation) due to its ability to preserve logical circuit-level parallelism. Within this context, there are also many choices e.g. whether to implement a CNOT gate using extended ancilla patches as in Ref. (Litinski and von Oppen, 2018) or using long-range Bell states as in Ref. (Beverland et al., 2022a), whether to implement an S gate using teleportation as in Ref. (Beverland et al., 2022a) or by braiding twist defects as in Ref. (Brown et al., 2017), or in how to perform a patch rotation following a transversal Hadamard gate (see Refs. (Vuillot et al., 2019; Litinski, 2019a; Beverland et al., 2022a; Horsman et al., 2012; Fowler and Gidney, 2018) for different protocols); the optimal choice will be similarly interdependent with choices of layout and hardware constraints.
@@
These considerations have motivated the realization that the three mentioned layers of the stack are highly inter-dependent and their co-optimization is formidable. At the same time, because the fault-tolerant layer is expected to dramatically increase the resource overhead of quantum computing, any end-to-end resource estimation pipeline for quantum circuits ought to include reasonable choices for it. To summarize this section, we take note that choices must be made in (at least) the following places: (a) a logical gate set with which to express input circuits (e.g. Clifford+T or Pauli product rotations), (b) a lattice surgery-native logical instruction set to translate input circuits to (e.g. the allowed set of Pauli product measurements), (c) an abstract layout specifying regions on the hardware dedicated to storing logical qubits, routing long-range instructions, and generating resource states (d) a protocol for implementing the lattice surgery-native logical instruction set using the layout (including methods for scheduling and routing these instructions as well as generating and managing any required resource states), and (e) a protocol for further compiling routed and scheduled lattice surgery instructions into native hardware gates. As we are convinced that extensible compilation and resource estimation platforms would greatly aid the community in optimizing resources within the fault-tolerant layer and culling the spectrum of propositions for surface code compilation strategies, we present one set of choices in this paper and expand the functionality of the LSC to suit.